OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)? y = −3x − 2 y = −3x + 2 y = 3x + 2 y = 3x − 2

OpenStudy (anonymous):

The slope is (10-4)/(4-2) = 6/3=3 So now you have 4=3(2) + b b = -2 The equation is y = 3x-2.

OpenStudy (anonymous):

Sorry I meant to type 6/2 = 3

OpenStudy (anonymous):

Thank you! I solved it already though but thanks for the help!